OpenStudy (anonymous):

Solution to 3/2 - 1/x+5 = -2/3x+3

OpenStudy (anonymous):

\[\frac{3}{2}-\frac{1}{x+5}=\frac{-2}{3x+3}\]?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

this is going to be a drag i would start with \[\frac{3}{2}=\frac{1}{x+5}-\frac{2}{3x+3}\] and then subtract on the right

OpenStudy (anonymous):

Subtract on the right how? I'm sorry

OpenStudy (anonymous):

/Combine the two expressions

OpenStudy (anonymous):

?

OpenStudy (anonymous):

\[\frac{1}{x+5}-\frac{2}{3x+3}=\frac{3x+3-2(x+5)}{(x+5)(3x+3)}\] then clean up the numerator

OpenStudy (anonymous):

The x+5's would "cancel" and is that it? There are also 3x+3's too, but one isn't held in parenthesis, so is this left alone?

OpenStudy (anonymous):

nothing cancels forget cancelling

OpenStudy (anonymous):

the numerator is \(3x+3-2x-10=x-7\) so the whole thing is \[\frac{x-7}{(x+5)(3x+3)}\]

OpenStudy (anonymous):

My end goal is to solve for x, so how would I move to the next step from here?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

"cross multiply" and get \[\frac{x-7}{(x+5)(3x+3)}=\frac{3}{2}\iff 2(x-7)=3(x+5)(3x+3) \]

OpenStudy (anonymous):

then you have an annoying quadratic equation to solve are you sure about the question ?

OpenStudy (anonymous):

Yea, I believe so, My question is "What is the solution to the rational expression.." and my options for the answer are x=..., with on eof the choices being no solution

OpenStudy (anonymous):

ok then if it was me, i would check the answers to see which one works

OpenStudy (anonymous):

Great idea!! Thank you :) I'll just plug in the values, If i need help, I'll just call you back. Thanks again :)

OpenStudy (anonymous):

k good luck i have a feeling either i made a mistake or maybe there was a typo in the question the solutions i got are strange maybe you could tell me what the possibilities are

OpenStudy (anonymous):

the answer options are x=-5, x=5, x=-5/2 and no solution

OpenStudy (anonymous):

then there was a typo in the question \[\frac{3}{2}-\frac{1}{x+5}=\frac{-2}{3x+3}\] does have solutions but none of those are you sure that is the exact question?

OpenStudy (anonymous):

Ahah! I've found it :p you're correct; the bottom portion of what is "-2/3x+3" is actually supposed to be 2x+10

OpenStudy (anonymous):

lordamercy!

OpenStudy (anonymous):

HAHA! Love your spelling on that, very visual lol

OpenStudy (anonymous):

then\[\frac{3}{2}-\frac{1}{x+5}=\frac{-2}{2x+10}\] is a ton easier, because the common denominator on the will be \(2x+10\) but after all this work i will tell you that the answer is "no solution"!!

OpenStudy (anonymous):

Alright, now that's cleared up, can you help me? haha

OpenStudy (anonymous):

you want to work it through again step by step or just go with the answer? i can go either way

OpenStudy (anonymous):

Yes the answer is fine :)

OpenStudy (anonymous):

Sorry I put you through all of that, lol, "lordamercy"

OpenStudy (anonymous):

kk good i am tired of this problem go with "no solution"

OpenStudy (anonymous):

LOL thanks again! you've been so much help tonight :D

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

btw a great "lordamercy" can be found here video is bad, but the lordamercy is excellent https://www.youtube.com/watch?v=lx_kwGttI2o

OpenStudy (anonymous):

Well that lordamercy is mighty fine if I say so myself, thank you for blessing me with that video, it's excellent

OpenStudy (anonymous):

and LOL to your comment hahaha

OpenStudy (anonymous):

big barrington levy fan bounty killer not so much

OpenStudy (anonymous):

"video is bad, but the lordamercy is excellent" what a perfect concoction of words haha

OpenStudy (anonymous):

I don't know of barrington, I'll have to look into him more. :) I like the reggae sound, which is what it reminds me of

OpenStudy (anonymous):

definitely reggae

OpenStudy (anonymous):

Lol, well I'm going to post my next question so I can get some friggin rest afterwards, my eyes are killing me!

OpenStudy (anonymous):

can't wait